OpenStudy (anonymous):

For the functions f(x) = 2x + 2 and g(x) = 7x + 1, which composition produces the greatest output? Both compositions produce the same output. Neither composition produces an output. f(g(x)) produces the greatest output. g(f(x)) produces the greatest output.

OpenStudy (anonymous):

someone help me please!!!!!!!

OpenStudy (anonymous):

yes king bruuh plz help me cause i like how you explain

OpenStudy (anonymous):

 I hope this help for doing 2x+2 but you

OpenStudy (anonymous):

but you'll have to fill out the rest of the chart if you want

OpenStudy (anonymous):



OpenStudy (anonymous):

But I think its g(X)= 7x+1 cause you pick any number for X plug it in and do the math you'll get a higher number than you would if you were to plug in a number the the f(X)= 2x+2

OpenStudy (anonymous):

so then its C

Vocaloid (vocaloid):

actually, I think you may be misunderstanding the problem

Vocaloid (vocaloid):

The question is not asking to compare f(x) and g(x), it's asking to compare f(g(x)) and g(f(x))

Vocaloid (vocaloid):

"composition" means to plug one function into the other function

OpenStudy (anonymous):

but i think its C

OpenStudy (anonymous):

W.e I was taught the way I sorta explained it. But @Papi_Chulo305 you can listen to the others just in case

Vocaloid (vocaloid):

what is f(g(x))?

OpenStudy (anonymous):

I also think it C

Vocaloid (vocaloid):

Hm, I don't think so f(g(x)) = 2(7x+1) + 2 = 14x + 2 + 2 = 14x + 4 g(f(x)) = 7(2x+2) + 1 = 14x + 14 + 1 = 14x + 15 g(f(x)) is bigger so I believe the answer is D

OpenStudy (anonymous):

THX

Vocaloid (vocaloid):

but do you understand what I did?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

thats what i meant

Vocaloid (vocaloid):

Oh, I see, sorry then D: